# LLM Personas as a Substitute for Field Experiments in Method Benchmarking

### v1

@@ Abstract
Field experiments (A/B tests) are often the most credible benchmark for methods in societal systems, but their cost and latency create a major bottleneck for iterative method development. LLM-based persona simulation offers a cheap synthetic alternative, yet it is unclear whether replacing humans with personas preserves the benchmark interface that adaptive methods optimize against. We prove an if-and-only-if characterization: when (i) methods observe only the aggregate outcome (aggregate-only observation) and (ii) evaluation depends only on the submitted artifact and not on the algorithm’s identity or provenance (algorithm-blind evaluation), swapping humans for personas is just panel change from the method’s point of view, indistinguishable from changing the evaluation population (e.g., New York to Jakarta). Furthermore, we move from validity to usefulness: we define an information-theoretic discriminability of the induced aggregate channel and show that making persona benchmarking as decision-relevant as a field experiment is fundamentally a sample-size question, yielding explicit bounds on the number of independent persona evaluations required to reliably distinguish meaningfully different methods at a chosen resolution.
@@ 1 Introduction
One of the recurring lessons from machine learning is that improvements in methods accelerate dramatically when we can iterate quickly against cheap and reliable evaluation benchmarks (blum2015ladder; zaharia2018accelerating; liao2021we; d2022underspecification; bommasani2023holistic; miller2024adding; abbas2025developing). The literature advances through an iterative process in which researchers propose a variant, test it against a benchmark, inspect failures, and refine. As a result, the limiting factor in progress is often the latency and reliability of the evaluation feedback loop rather than the availability of new ideas (xin2018developers).
@@
In many societal applications, such as pricing and matching policies, ad allocation, marketplace design, and behavioral interventions, the closest analogue to a benchmark is a field experiment (harrison2004field; duflo2017economist), also known as A/B testing (kohavi2020trustworthy). Here, competing methods can be compared by deploying them through estimating their causal effects on agreed-upon outcome metrics. Repeating such experiments across settings allows researchers to assess not only average performance but also robustness and behavior under adaptation.
@@
However, field experiments are costly to set up and slow to run (diamond1986laboratory; fisher1992arrangement; paluck2014field; samek2019advantages). Researchers typically must first collaborate with an organization, propose an intervention, negotiate approval across multiple stakeholders, and spend months implementing instrumentation and experimental infrastructure before any data are collected, and then wait to collect data until enough statistical power is gathered. Moreover, feasible experiments are often constrained by the organization’s interests, risk tolerance, and resource limitations, which restrict what methods can be tested and how frequently experiments can be repeated (bandiera2011field).
@@
Recently, LLM-based persona simulation has shown potential to serve as a cheaper synthetic field laboratory (toubia2025database; peng2025mega). In LLM-based persona simulation, a large language model is conditioned on explicit persona descriptions, such as survey responses that capture demographic attributes, preferences, roles, or beliefs. This is then used to generate individual-level evaluations or responses to proposed policies, content, or system outputs. These simulated micro-level responses can then be aggregated into a single evaluation signal. Because such simulations are inexpensive and fast to run at scale, they make it feasible to test and iterate on methods without field experimentation or A/B testing.
@@
However, whether LLM-based persona simulation can indeed reliably substitute for field experimentation or A/B testing for methodology testing remains largely unknown. This uncertainty is exacerbated by recent studies that provide negative results in applications beyond methodology testing, particularly when we ask causal questions to support external validity claims (gui2023challenge; li2025llm; peng2025mega; gui2025leveraging). Manipulating prompts intended to represent a treatment can inadvertently shift other latent aspects of the scenario, inducing confounding.
@@
    algorithm-blind evaluation (AB): For each method, the distribution of the returned score depends only on what was submitted and not on which training/optimization procedure produced.
@@
In addition to this identification result, we extend the identification discussion into a notion of usefulness: once (AO)+(AB) makes persona benchmarking identification-valid (i.e., “just panel change” on the method’s interface), the remaining question is whether the induced aggregate channel $Q_{\mathrm{pers}}(\cdot\mid w)$ is informative enough to distinguish and optimize meaningfully different methods. We formalize this via an information-based discriminability parameter (worst-case KL separation at a pre-specified resolution) and show it yields an explicit sample-complexity rule for how many independent persona evaluations (equivalently, what effective persona dataset/panel size) are required for reliable method comparison. In this sense, beyond enforcing (AO)+(AB), “persona quality” becomes a measurable budget question: is the persona panel large enough to resolve the improvements we care about?
The rest of the paper is organized as follows. Section 2 reviews related work. Section 3 formalizes the benchmarking setup and states the (AO) and (AB) conditions. Section 4 gives the main identification discussion, proving that if and only if (AO) and (AB) holds, persona vs. human evaluation is just panel change (JPC). Section 5 extends the identification discussion to a usefulness discussion that relates to sample complexity. The Appendix contains deferred proofs.
@@ 2 Related works
Aggregate-only observation (AO), a key condition for using persona simulation rather than field experiments for method benchmarking, has also appeared in the literature on benchmark gaming (blum2015ladder; hardt2017climbing; feldman2019advantages; biderman2024trenches) as a baseline for preventing leaderboard gaming via adaptive overfitting. In econometrics, treatments of randomized experiments formalize the evaluation protocol as an assignment-and-aggregation procedure, clarifying what is identified from the realized (often aggregated) outcomes (atheyimbens2017econometrics).
Another key condition, algorithm-blind evaluation (AB), has been experimentally motivated: people rate the same advice differently when it is labeled as AI rather than human/crowd-sourced (bogert2022human; osborne2025me); many large-scale evaluation platforms implement partial blinding (e.g., anonymous pairwise comparisons) precisely to reduce provenance effects (chiang2024chatbotarena). dominguezolmedo2024questioning shows that survey-style elicitation of LLMs exhibits strong ordering and labeling effects that can dominate conclusions. The econometrics literature has also shown that outcomes can shift when participants condition on contextual/provenance cues, motivating blinding and stable scoring protocols (dequidt2018demand; levitt2011hawthorne).
Related to quantifying a benchmark’s usefulness, madaan2024variance studied variance-focused analyses that quantify the extent to which benchmark scores fluctuate during evaluation. heineman2025signal discusses “signal vs. noise” perspectives and argue that benchmarks with higher signal-to-noise ratios are more reliable for model-selection decisions. Compared to these papers, we focus on proposing a benchmark-internal information measure (minimum KL separation) that directly yields sample-size scaling for reliable method comparison on the induced aggregate channel, clarifying how large a persona dataset is practically enough.
@@ 3 Setup: Benchmarking experimentation
3.1 knobs, artifacts, and algorithms
Let $\Theta$ denote the space of method configuration knobs: all controllable degrees of freedom that specify a system or a procedure (e.g., model weights, prompts/context, hyperparameters, decoding rules, tool policies, memory policies, data curation choices, or post-processing rules). A method configuration is $\theta\in\Theta$.
Deploying $\theta$ yields an evaluable artifact
$w(\theta)\in\mathcal{W},$ (1)
where $\mathcal{W}$ is the artifact space. Intuitively, an artifact $w(\theta)$ is what you hand to the benchmark, i.e., what evaluator queries. Possible choice of the artifact space $\mathcal{W}$ encompass: (i) a single output for a fixed input, (ii) an output distribution (stochastic method), (iii) a full interaction method mapping contexts to actions/outputs, (iv) a rollout distribution of an agent interacting with tools or environments.
Algorithm as an interactive learner.
We model algorithm benchmarking as a repeated “submit–observe” loop. At each round $t=1,2,\dots,T$, the algorithm chooses a configuration $\theta_{t}\in\Theta$ (equivalently an artifact $w_{t}:=w(\theta_{t})\in\mathcal{W}$), submits $w_{t}$ to an evaluator, and receives a feedback observation $o_{t}$ taking values in some observation space $\mathcal{O}$. Formally, an algorithm $\mathcal{A}$ is modeled as a (possibly randomized and adaptive) procedure that induces a decision kernel $\pi_{t}$ over configurations at each round $t$, such that
@@
where $H_{t-1}$ is the algorithm’s observable history before round $t$, i.e.,
@@
Note that such a definition of an algorithm contains:
  • Offline alignment (e.g., DPO (rafailov2023direct), SFT (wei2022finetuned) ) as the special case where $\pi_{t}$ does not depend on $H_{t-1}$ (or where $T=1$).
  • Online alignment (e.g., RLHF (ouyang2022instructgpt)) as the case where $\pi_{t}$ adapts to past feedback $o_{1:t-1}$.
  • Hyperparameter tuning / AutoML / architecture search (he2021automl; ren2021comprehensive) as black-box optimization over $\Theta$ using benchmark scores, where $\theta_{t}$ encodes a full training-and-deployment recipe, and $\pi_{t}$ implements a sequential search procedure.
  • Prompt and system configuration search (pryzant2023automatic; kang2025bayesian) as the case where $\Theta$ indexes prompts, system messages, tool-use policies, retrieval and memory settings, and post-processing rules.
  • Data-centric training recipe search (zha2025data) as the case where $\Theta$ includes dataset construction and curation choices (filtering, reweighting, mixing, synthetic-data generation policies).
@@ 3.2 Evaluation as panel $\times$ instrument $\times$ aggregation
We model an evaluation as a two-stage process: many micro-level judgments are first produced, and these are then compressed into a single aggregate feedback signal that the method actually observes. Under this model, an evaluation setup (or simply an evaluator) is fully specified by a tuple
@@
where $\mathcal{O}$ is the method’s feedback space. Typical examples include the mean score, a majority vote, or a pass/fail indicator. Putting these pieces together, a single evaluation call on artifact $w$ returns the aggregate
@@
Although evaluation involves panel members and micro-responses, the method never observes these internal details. What it sees is only the induced distribution of the aggregate feedback $o\in\mathcal{O}$. The tuple $(P,I,\Gamma,L)$ therefore defines a Markov kernel on $\mathcal{O}$:
@@
Although these two pipelines differ internally (humans vs. personas; human judgments vs. LLM judgments), the algorithm only interacts with each benchmark through the induced distribution of the aggregate feedback. In other words, for the method, the relevant objects are precisely the two reduced-form kernels $Q_{\mathrm{hum}}(\cdot\mid w)$ and $Q_{\mathrm{pers}}(\cdot\mid w)$.
@@ 3.4 Key assumptions
In asking the key question in this paper, “when can we treat persona evaluation as a clean benchmark interface for comparing algorithms?”, we need to discuss two “benchmark hygiene” conditions that clarify what information the method does (and does not) get access to, and whether the benchmark behaves like a well-defined environment independent of who is playing. They are not substantive modeling assumptions about humans or LLMs; they are assumptions about what information the benchmark reveals and how the evaluator behaves as an environment.
@@
At each round $t$, the algorithm observes only the aggregate feedback $o_{t}\in\mathcal{O}$. It does not observe the micro-level tuple $(p_{1},\dots,p_{L},Z_{1},\dots,Z_{L})$, any panel identifiers, or any additional side-channel information beyond $o_{t}$.
Intuitively, (AO) says the method sees exactly what a standard leaderboard would show: one score (or label) per submission. This prevents algorithm’s gaming behavior that relies on recognizing individual panelists/personas or exploiting micro-level structure that would be invisible in the intended benchmark interface.
Before discussing the next assumption, we define a probability measure $\mathbb{P}^{\mathcal{A}}$: for a fixed benchmark implementation, running an algorithm $\mathcal{A}$ induces a probability measure $\mathbb{P}^{\mathcal{A}}$ over the interaction transcript $(w_{1},o_{1},w_{2},o_{2},\dots)$. Under (AO), we can define the algorithm’s interaction transcript after t1t-1 rounds as
@@
Assumption 2 (Algorithm-blind evaluation (AB)).
There exists a Markov kernel $Q(\cdot\mid w)$ on $\mathcal{O}$ such that for every algorithm $\mathcal{A}$, every round $t$, and every measurable $A\subseteq\mathcal{O}$,
@@
Intuitively, (AB) is the minimal condition for calling this evaluation setup a benchmark environment at all: the evaluator should not care about the identity of the training procedure or other metadata, and care only about what was submitted (the artifact). In other words, the benchmark interaction is fully summarized by the reduced-form kernel $Q(\cdot\mid w)$, which is fixed across algorithms.
@@ 4 Identification: When Is Persona Benchmarking “Just Panel Change”?
We first formalize what it means to “only change the evaluation population” in the panel–instrument–aggregation model (a literal panel change), and then define the corresponding interface-level notion that is relevant for adaptive benchmarking methods (just panel change, JPC). We then show that for human vs. persona benchmarking, JPC holds if and only if two benchmark-hygiene conditions hold: aggregate-only observation (AO) and algorithm-blind evaluation (AB).
This yields an identification result: under (AO)+(AB), swapping humans for personas is indistinguishable on the method interface from an ordinary change of evaluation population (even though the internal micro-instrument may differ); conversely, if either condition fails, the swap can change the interface in ways that go beyond “panel change.”

### v2

@@ Abstract
Field experiments (A/B tests) are often the most credible benchmark for methods (algorithms) in societal systems, but their cost and latency bottleneck rapid methodological progress. LLM-based persona simulation offers a cheap synthetic alternative, yet it is unclear whether replacing humans with personas preserves the benchmark interface that adaptive methods optimize against. We prove an if-and-only-if characterization: when (i) methods observe only the aggregate outcome (aggregate-only observation) and (ii) evaluation depends only on the submitted artifact and not on the method’s identity or provenance (method-blind evaluation), swapping humans for personas is just panel change from the method’s point of view, indistinguishable from changing the evaluation population (e.g., New York to Jakarta). Furthermore, we move from validity to usefulness: we define an information-theoretic discriminability of the induced aggregate channel and show that making persona benchmarking as decision-relevant as a field experiment is fundamentally a sample-size question, yielding explicit bounds on the number of independent persona evaluations required to reliably distinguish meaningfully different methods at a chosen resolution.
@@ 1 Introduction
One of the recurring lessons from machine learning is that improvements in methods accelerate dramatically when we can iterate quickly against cheap and reliable evaluation benchmarks (Blum and Hardt, 2015; Zaharia et al., 2018; Liao et al., 2021; D’Amour et al., 2022; Bommasani et al., 2023; Miller, 2024; Abbas et al., 2025). The literature advances through an iterative process in which researchers propose a variant, test it against a benchmark, inspect failures, and refine. As a result, the limiting factor in progress is often the latency and reliability of the evaluation feedback loop rather than the availability of new ideas (Xin et al., 2018).
@@
In many societal applications, such as pricing and matching policies, ad allocation, marketplace design, and behavioral interventions, the closest analogue to a benchmark is a field experiment (Harrison and List, 2004; Duflo, 2017), also known as A/B testing (Kohavi et al., 2020). Here, competing methods can be compared by deploying them through estimating their causal effects on agreed-upon outcome metrics. Repeating such experiments across settings allows researchers to assess not only average performance but also robustness and behavior under adaptation.
@@
However, field experiments are costly to set up and slow to run (Diamond, 1986; Fisher, 1992; Paluck and Cialdini, 2014; Samek, 2019). Researchers typically must first collaborate with an organization, propose an intervention, negotiate approval across multiple stakeholders, and spend months implementing instrumentation and experimental infrastructure before any data are collected, and then wait to collect data until enough statistical power is gathered. Moreover, feasible experiments are often constrained by the organization’s interests, risk tolerance, and resource limitations, which restrict what methods can be tested and how frequently experiments can be repeated (Bandiera et al., 2011).
@@
Recently, LLM-based persona simulation has shown potential to serve as a cheaper synthetic field laboratory (Toubia et al., 2025; Peng et al., 2025). In LLM-based persona simulation, a large language model is conditioned on explicit persona descriptions, such as survey responses that capture demographic attributes, preferences, roles, or beliefs. This is then used to generate individual-level evaluations or responses to proposed policies, content, or system outputs. These simulated micro-level responses can then be aggregated into a single evaluation signal. Because such simulations are inexpensive and fast to run at scale, they make it feasible to test and iterate on methods without field experimentation or A/B testing.
@@
However, whether LLM-based persona simulation can indeed reliably substitute for field experimentation or A/B testing for methodology testing remains largely unknown. This uncertainty is exacerbated by recent studies that provide negative results in applications beyond methodology testing, particularly when we ask causal questions to support external validity claims (Gui and Toubia, 2023; Li et al., 2025; Peng et al., 2025; Gui and Kim, 2025). Manipulating prompts intended to represent a treatment can inadvertently shift other latent aspects of the scenario, inducing confounding.
@@
    method-blind evaluation (MB): For each method, the distribution of the returned score depends only on what was submitted and not on which training/optimization procedure produced.
@@
In addition to this identification result, we extend the identification discussion into a notion of usefulness: once (AO)+(MB) makes persona benchmarking identification-valid (i.e., “just panel change” on the method’s interface), the remaining question is whether the induced aggregate channel $Q_{\mathrm{pers}}(\cdot\mid w)$ is informative enough to distinguish and optimize meaningfully different methods. We formalize this via an information-based discriminability parameter (worst-case KL separation at a pre-specified resolution) and show it yields an explicit sample-complexity rule for how many independent persona evaluations (equivalently, what effective persona dataset/panel size) are required for reliable method comparison. In this sense, beyond enforcing (AO)+(MB), “persona quality” becomes a measurable budget question: is the persona panel large enough to resolve the improvements we care about?
The rest of the paper is organized as follows. Section 2 reviews related work. Section 3 formalizes the benchmarking setup and states the (AO) and (MB) conditions. Section 4 gives the main identification discussion, proving that if and only if (AO) and (MB) holds, persona vs. human evaluation is just panel change (JPC). Section 5 extends the identification discussion to a usefulness discussion that relates to sample complexity. The Appendix contains deferred proofs.
@@ 2 Related works
Aggregate-only observation (AO), a key condition for using persona simulation rather than field experiments for method benchmarking, has also appeared in the literature on benchmark gaming (Blum and Hardt, 2015; Hardt, 2017; Feldman et al., 2019; Biderman et al., 2024) as a baseline for preventing leaderboard gaming via adaptive overfitting. In econometrics, treatments of randomized experiments formalize the evaluation protocol as an assignment-and-aggregation procedure, clarifying what is identified from the realized (often aggregated) outcomes (Athey and Imbens, 2017).
Another key condition, method-blind evaluation (MB), has been experimentally motivated: people rate the same advice differently when it is labeled as AI rather than human/crowd-sourced (Bogert et al., 2022; Osborne and Bailey, 2025); many large-scale evaluation platforms implement partial blinding (e.g., anonymous pairwise comparisons) precisely to reduce provenance effects (Chiang et al., 2024). Dominguez-Olmedo et al. (2024) shows that survey-style elicitation of LLMs exhibits strong ordering and labeling effects that can dominate conclusions. The econometrics literature has also shown that outcomes can shift when participants condition on contextual/provenance cues, motivating blinding and stable scoring protocols (de Quidt et al., 2018; Levitt and List, 2011).
Related to quantifying a benchmark’s usefulness, Madaan et al. (2024) studied variance-focused analyses that quantify the extent to which benchmark scores fluctuate during evaluation. Heineman et al. (2025) discusses “signal vs. noise” perspectives and argue that benchmarks with higher signal-to-noise ratios are more reliable for model-selection decisions. Compared to these papers, we focus on proposing a benchmark-internal information measure (minimum KL separation) that directly yields sample-size scaling for reliable method comparison on the induced aggregate channel, clarifying how large a persona dataset is practically enough.
@@ 3 Setup: Benchmarking experimentation
3.1 Configurations and artifacts.
Let $\Theta$ denote the space of tunable configuration variables; this includes all controllable degrees of freedom that specify a system or a procedure (e.g., model weights, prompts/context, hyperparameters, decoding rules, tool policies, memory policies, data curation choices, or post-processing rules). A single candidate configuration is a choice $\theta\in\Theta$. A submitted artifact is what you hand to the benchmark, i.e., the externally visible object that the benchmark evaluates. We model this via an artifact map $g:\Theta\to\mathcal{W}$ and write
$w=g(\theta)\in\mathcal{W}.$
Possible choice of the artifact space $\mathcal{W}$ encompass: (i) a single output for a fixed input, (ii) an output distribution (stochastic method), (iii) a full interaction method mapping contexts to actions/outputs, (iv) a rollout distribution of an agent interacting with tools or environments.
Method as a configuration optimizer.
We model method benchmarking as a “submit–observe” loop. A method (algorithm) may be either non-adaptive (a single-shot submission, or a fixed distribution over submissions) or adaptive (updating submissions based on past benchmark feedback). Concretely, at each round $t=1,2,\dots,T$, the method chooses a configuration $\theta_{t}\in\Theta$ (equivalently an artifact $w_{t}:=g(\theta_{t})\in\mathcal{W}$), submits $w_{t}$ to an evaluator, and receives a feedback observation $o_{t}$ taking values in some observation space $\mathcal{O}$. Formally, a method (or method) $\mathcal{A}$ is modeled as a (possibly randomized) procedure that induces a decision kernel $\pi_{t}$ over configurations at each round $t$, such that
@@
where $H_{t-1}$ is the method’s observable history before round $t$, i.e.,
@@
Note that such a definition of an method contains:
  • Offline alignment (e.g., DPO (Rafailov et al., 2023), SFT (Wei et al., 2022) ) as the special case where $\pi_{t}$ does not depend on $H_{t-1}$ (or where $T=1$).
  • Online alignment (e.g., RLHF (Ouyang et al., 2022)) as the case where $\pi_{t}$ adapts to past feedback $o_{1:t-1}$.
  • Hyperparameter tuning / AutoML / architecture search (He et al., 2021; Ren et al., 2021) as black-box optimization over $\Theta$ using benchmark scores, where $\theta_{t}$ encodes a full training-and-deployment recipe, and $\pi_{t}$ implements a sequential search procedure.
  • Prompt and system configuration search (Pryzant et al., 2023; Kang and Yoganarasimhan, 2025) as the case where $\Theta$ indexes prompts, system messages, tool-use policies, retrieval and memory settings, and post-processing rules.
  • Data-centric training recipe search (Zha et al., 2025) as the case where $\Theta$ includes dataset construction and curation choices (filtering, reweighting, mixing, synthetic-data generation policies).
@@ 3.2 Evaluation as panel $\times$ instrument $\times$ aggregation
We model an evaluation as a two-stage process: many micro-level judgments are first produced, and these are then compressed into a single aggregate feedback signal that the method actually observes. Under this model, an evaluation setup (or simply an evaluator) is fully specified as
@@
where $\mathcal{O}$ is the feedback observation space. Typical examples include the mean score, a majority vote, or a pass/fail indicator. Putting these pieces together, a single evaluation call on artifact $w$ returns the aggregate
@@
Although evaluation involves panel members and micro-responses, the method never observes them. What it sees is only the induced distribution of the aggregate feedback $o\in\mathcal{O}$. The tuple $(P,I,\Gamma,L)$ therefore defines a Markov kernel on $\mathcal{O}$:
@@
Although these two pipelines differ internally (humans vs. personas; human judgments vs. LLM judgments), the method only interacts with each benchmark through the induced distribution of the aggregate feedback. In other words, for the method, the relevant objects are precisely the two reduced-form kernels $Q_{\mathrm{hum}}(\cdot\mid w)$ and $Q_{\mathrm{pers}}(\cdot\mid w)$.
@@ 3.4 Key assumptions
In asking the key question in this paper, “when can we treat persona evaluation as a clean benchmark interface for comparing methods?”, we need to discuss two “benchmark hygiene” conditions that clarify what information the method does (and does not) get access to, and whether the benchmark behaves like a well-defined environment independent of who is playing. They are not substantive modeling assumptions about humans or LLMs; they are assumptions about what information the benchmark reveals and how the evaluator behaves as an environment.
@@
At each round $t$, the method observes only the aggregate feedback $o_{t}\in\mathcal{O}$. It does not observe the micro-level tuple $(p_{1},\dots,p_{L},Z_{1},\dots,Z_{L})$, any panel identifiers, or any additional side-channel information beyond $o_{t}$.
Intuitively, (AO) says the method sees exactly what a standard leaderboard would show: one score (or label) per submission. This prevents method’s gaming behavior that relies on recognizing individual panelists/personas or exploiting micro-level structure that would be invisible in the intended benchmark interface.
Before discussing the next assumption, we define a probability measure $\mathbb{P}^{\mathcal{A}}$: for a fixed benchmark implementation, running an method $\mathcal{A}$ induces a probability measure $\mathbb{P}^{\mathcal{A}}$ over the interaction transcript $(w_{1},o_{1},w_{2},o_{2},\dots)$. Under (AO), we can define the method’s interaction transcript after t1t-1 rounds as
@@
Assumption 2 (Method-blind evaluation (MB)).
There exists a Markov kernel $Q(\cdot\mid w)$ on $\mathcal{O}$ such that for every method $\mathcal{A}$, every round $t$, and every measurable $A\subseteq\mathcal{O}$,
@@
Intuitively, (MB) is the minimal condition for calling this evaluation setup a benchmark environment at all: the evaluator should not care about the identity of the training procedure or other metadata, and care only about what was submitted (the artifact). In other words, the benchmark interaction is fully summarized by the reduced-form kernel $Q(\cdot\mid w)$, which is fixed across methods.
@@ 4 Identification: When Is Persona Benchmarking “Just Panel Change”?
We first formalize what it means to “only change the evaluation population” in the panel–instrument–aggregation model (a literal panel change), and then define the corresponding interface-level notion that is relevant for adaptive benchmarking methods (just panel change, JPC). We then show that for human vs. persona benchmarking, JPC holds if and only if two benchmark-hygiene conditions hold: aggregate-only observation (AO) and method-blind evaluation (MB).
This yields an identification result: under (AO)+(MB), swapping humans for personas is indistinguishable on the method interface from an ordinary change of evaluation population (even though the internal micro-instrument may differ); conversely, if either condition fails, the swap can change the interface in ways that go beyond “panel change.”